A stone is thrown upward from the top of a building at an angle of 30.0° to the


horizontal and with an initial speed of 20.0 m/s. The point of release is 45.0 m above the


ground.


a. How long does it take to hit the ground?


b. Find the stone’s speed impact.


c. Find the horizontal range of the stone.

1
Expert's answer
2022-02-11T10:17:14-0500

Explanations & Calculations


  • This is a comprehensive questionnaire to understand the projectile motion under gravity.

a)

  • Apply "s=ut +\frac{1}{2}at^2" upwards on the stone for its motion starting at the throw until the impact. So at the impact, it would have travelled a displacement of "\small -45.0\,m". And the equation to solve would be like

"\qquad\qquad
\begin{aligned}
\small \uparrow\qquad -45&=\small (20.0\sin30)t+\frac{1}{2}(-9.8)t^2
\end{aligned}"

  • You can solve the quadratic equation formed here and find the time of flight. Neglect any negative values....[4.21s]


b)

  • Apply energy conservation with respect to the impact level.

"\qquad\qquad
\begin{aligned}
\small mgh+\frac{1}{2}mv_1^2&=\small mgh+\frac{1}{2}mv_2^2\\
\small mg(45)+\frac{1}{2}m(20)^2&=\small 0+\frac{1}{2}.m.v_2^2\\
\small 9.8\times45+\frac{1}{2}.(20)^2&=\small \frac{v_2^2}{2}
\end{aligned}"

  • Solve for "\small v_2" .


c)

  • The all-time constant horizontal velocity is "\small v\cos\theta=20\cos30"
  • Then the range can simply be calculated by

"\qquad\qquad
\begin{aligned}
\small \to \qquad s&=\small ut\\
&=\small 20\cos30\times \text{time of flight}
\end{aligned}"